OpenStudy (anonymous):

Allie has a birthday present for her brother. The length of the box that the present comes in is two times longer than the width of the box. The height of the box is 5 inches and its volume is 160 cubic inches. What is the length and width of the box?

OpenStudy (anonymous):

Equation: Volume=W*L*H knowns: H=5 inches Volume=160 cubic inches 2*L=W Attempt at solution: At this point we want to substitute height, volume and 2*L for W and solve for L. After that you plug back L to the original equation and solve for W

OpenStudy (anonymous):

thanks!